# Rethinking Multilingual Continual Pretraining: Data Mixing for Adapting LLMs Across Languages and Resources

### v1

@@ Abstract
Large Language Models (LLMs) exhibit significant disparities in performance across languages, primarily benefiting high-resource languages while marginalizing underrepresented ones. Continual Pretraining (CPT) has emerged as a promising approach to address this imbalance, although the relative effectiveness of monolingual, bilingual, and code-augmented data strategies remains unclear. This study systematically evaluates 36 CPT configurations involving three multilingual base models, across 30+ languages categorized as altruistic, selfish, and stagnant, spanning various resource levels. Our findings reveal three major insights: (1) Bilingual CPT improves multilingual classification but often causes language mixing issues during generation. (2) Including programming code data during CPT consistently enhances multilingual classification accuracy, particularly benefiting low-resource languages, but introduces a trade-off by slightly degrading generation quality. (3) Contrary to prior work, we observe substantial deviations from language classifications according to their impact on cross-lingual transfer: Languages classified as altruistic often negatively affect related languages, selfish languages show conditional and configuration-dependent behavior, and stagnant languages demonstrate surprising adaptability under certain CPT conditions. These nuanced interactions emphasize the complexity of multilingual representation learning, underscoring the importance of systematic studies on generalizable language classification to inform future multilingual CPT strategies.
@@ 1 Introduction
Large Language Models (LLMs), built upon the Transformer architecture (Vaswani et al., 2017), have achieved remarkable progress in tasks such as machine translation, text classification, and generative dialogue. Despite these advances, their performance remains highly uneven across languages, favoring high-resource languages and marginalizing underrepresented ones (Li et al., 2024). This imbalance deepens the digital language divide and limits the inclusivity of NLP technologies.
@@
In addition to textual data in natural languages, a growing practice in LLM training is to incorporate programming code as an additional source of information. Previous research indicates that incorporating code enhances reasoning capabilities and improves the ability to handle structured information (Petty et al., 2024; Aryabumi et al., 2024), but its role in multilingual context remains underexplored.
There is a critical gap in understanding how language characteristics interact with CPT strategies. A recent classification proposed by Yuan et al. (2024) categorizes languages as altruistic, selfish, and stagnant based on their cross-lingual transfer patterns. However, this classification has only been validated in narrow experimental settings using English-centric bilingual data, leaving open questions about its generalizability to: (1) non-English language pairs, (2) code-augmented training regimes, and (3) models with varying pretraining corpora and architectures.
@@ 2.1 Language Selection
For each category, we select 1 high-resource language (except for the stagnant category for which no high-resource language is available in the dataset we use), 2 medium-resource languages, and 2 low-resource languages to ensure a balanced representation across different resource levels. The classification of languages into high-, medium-, and low-resource categories is determined by analyzing the data distribution of the Lego-MT dataset (Yuan et al., 2023), which serves as the basis for our setup. Specifically, we calculate the total token count for each language. Languages are then categorized as follows: high-resource languages exceed 1 billion tokens, medium-resource languages range between 10 million and 1 billion tokens, and low-resource languages fall below 10 million tokens. These languages serve as the training languages in our CPT experiments. The selected training languages, along with their corresponding category and resource level, are summarized in the first three columns of Table 1.
@@ 2.2 Pretraining Data
We finally select data for 15 languages across our three categories: altruistic (6 languages, ~92M tokens), selfish (6 languages, ~100M tokens), and stagnant (5 languages including English, ~87M tokens), resulting in a total of approximately 279 million tokens.
@@
For training configurations that include code, we maintain a 2:1 ratio between textual (monolingual/bilingual) and code data, with code comprising about 33% of the total tokens. This aligns with prior work (Aryabumi et al., 2024), which recommends a 25% code proportion (text:code ~3:1) for balancing language and code performance, noting that 33% remains reasonable for enhancing reasoning tasks. We sample the code dataset down to 50 million tokens, matching the 100 million tokens of textual data.
@@ 2.4 CPT Configurations
[…]
@@
[…]
@@
Base Model Category Training Data
Mono Bi Mono+Code Bi+Code
Selfish L3-Mono-Sel L3-Bi-Sel L3-Mono+Code-Sel L3-Bi+Code-Sel
Selfish L2-Mono-Sel L2-Bi-Sel L2-Mono+Code-Sel L2-Bi+Code-Sel
Selfish V7-Mono-Sel V7-Bi-Sel V7-Mono+Code-Sel V7-Bi+Code-Sel
Table 2: Continual pretraining configurations with structured naming.
@@ 3.2.2 Comparative Analysis in Classification Tasks
Mid-resource languages show mixed trends. Llama-3.1-8B maintains near-baseline performance with bilingual CPT (64.05% vs baseline 65.85%, -2.7%), but monolingual CPT degrades significantly (52.53%, -20.2%). Llama-2-7B struggles across both configurations, with bilingual CPT reducing accuracy to 24.26% (vs baseline 20.59%, +17.8%) and monolingual CPT performing slightly better (26.80%, +30.2%). Viking-7B uniquely benefits from bilingual CPT (29.74% vs baseline 23.94%, +24.2%), while monolingual CPT underperforms (24.02%, +0.3%). This indicates that bilingual CPT can stabilize mid-resource language performance for certain models (e.g., Viking-7B and Llama-2-7B). However, monolingual CPT risks overfitting to limited in-language data, particularly for models with weaker pretraining (e.g., Llama-3.1-8B).
@@
Low-resource languages exhibit divergent patterns. Llama-3.1-8B improves with bilingual CPT (62.91% vs baseline 56.86%, +10.6%) but declines with monolingual CPT (52.04%, -8.5%). Llama-2-7B degrades significantly with bilingual CPT (20.84%, +19.8%) and shows minimal gains with monolingual CPT (24.51%, +40.9%). Viking-7B benefits substantially from bilingual CPT (29.25%, +66.5%), while monolingual CPT slightly underperforms (21.33%, +21.4%). This highlights that bilingual CPT can enhance low-resource language performance for models with compatible pretraining (e.g., Viking-7B and Llama-3.1-8B).
@@ 4 Conclusion
First, while bilingual CPT enhances classification accuracy for mid- and low-resource languages, it introduces language mixing during generation, limiting its utility for translation tasks. Second, code integration during CPT acts as a scaffold for low-resource language understanding but introduces task-dependent trade-offs, improving classification while slightly degrading generation quality. Third, we demonstrate that language classifications based on cross-lingual transfer patterns (altruistic, selfish, stagnant) fail to generalize under varying CPT strategies.
@@
All resources are available at https://mala-lm.github.io/MixCPT.html.
@@ A.1 Related Work
Continual Pretraining
Continual pretraining has emerged as a pivotal technique for adapting LLMs to new domains or languages while retaining previously acquired knowledge (Yıldız et al., 2024). This approach has demonstrated significant benefits across diverse domains, including cybersecurity (Yu et al., 2025), finance(Hirano & Imajo, 2024), and law (Niyogi & Bhattacharya, 2024). In the context of language adaptation, researchers have successfully leveraged continual pretraining to enhance performance on low- and medium-resource languages. For instance, Ji et al. (2024a); Lu et al. (2024) extended the capabilities of open-weight LLMs by pretraining them on multilingual datasets encompassing hundreds of languages. Similarly, Fujii et al. (2024) significantly improved Japanese language proficiency by continually pretraining LLama-2 (Touvron et al., 2023) on a large-scale Japanese web corpus. In another study, Vo et al. (2024) achieved notable advancements in Korean language processing by utilizing 9.7 billion tokens for continual pretraining.
@@
A.2 Additional Results on Bilingual CPT with Code Data
@@ A.3 SIB-200 Accuracy
A.4 FLORES-200 BLEU Scores
@@
A.5 Language Mixing Examples

### v2

@@ Abstract
Large Language Models (LLMs) exhibit significant disparities in performance across languages, primarily benefiting high-resource languages while marginalizing underrepresented ones. Continual Pretraining (CPT) has emerged as a promising approach to address this imbalance, although the relative effectiveness of monolingual, bilingual, and code-augmented data strategies remains unclear. This study systematically evaluates 36 CPT configurations involving three multilingual base models, across 30+ languages categorized as altruistic, selfish, and stagnant, spanning various resource levels. Our findings reveal three major insights: (1) Bilingual CPT improves multilingual classification but often causes language mixing issues during generation. (2) Including programming code data during CPT consistently enhances multilingual classification accuracy and language modeling capabilities, particularly benefiting low-resource languages, but introduces a trade-off by slightly degrading generation quality. (3) Contrary to prior work, we observe substantial deviations from language classifications according to their impact on cross-lingual transfer: Languages classified as altruistic often negatively affect related languages, selfish languages show conditional and configuration-dependent behavior, and stagnant languages demonstrate surprising adaptability under certain CPT conditions. These nuanced interactions emphasize the complexity of multilingual representation learning, underscoring the importance of systematic studies on generalizable language classification to inform future multilingual CPT strategies.
@@ 1 Introduction
Large Language Models (LLMs), built upon the Transformer architecture (Vaswani et al., 2017), have achieved remarkable progress in tasks such as machine translation, text classification, and generative dialogue. Despite these advances, their performance remains highly uneven across languages, favoring high-resource languages and marginalizing underrepresented ones (Li et al., 2024; Joshi et al., 2020; Choudhury, 2023). This imbalance deepens the digital language divide and limits the inclusivity of NLP technologies.
@@
In addition to textual data in natural languages, a growing practice in LLM training is to incorporate programming code. For multilingual models, the inclusion of code, with its distinct logical syntax and lower ambiguity compared to natural language, can diversify the training data. We hypothesize that this diversity compels the model to learn more general and robust representations, moving beyond language-specific surface patterns; thus improving both the model’s predictive ability and its performance on downstream tasks.
There is a critical gap in understanding how language characteristics interact with CPT strategies. A recent classification proposed by Yuan et al. (2024) categorizes languages as altruistic, selfish, and stagnant based on their cross-lingual transfer patterns. This framework classifies languages based on their impact in multilingual training: altruistic languages are hypothesized to improve performance on related languages, selfish languages primarily benefit themselves, and stagnant languages offer little to no benefit to any language, including their own. However, this classification has only been validated in narrow experimental settings using English-centric bilingual data, leaving open questions about its generalizability to: (1) non-English language pairs, (2) code-augmented training regimes, and (3) models with varying pretraining corpora and architectures.
@@ 2.1 Language Selection
For each category, we select 1 high-resource language (except for the stagnant category, for which no high-resource language is available in the dataset we use), 2 medium-resource languages, and 2 low-resource languages to ensure a balanced representation across different resource levels. The classification of languages into high-, medium-, and low-resource categories is determined by analyzing the data distribution of the Lego-MT dataset (Yuan et al., 2023), which serves as the basis for our setup. Specifically, we calculate the total token count for each language. Languages are then categorized as follows: high-resource languages exceed 1 billion tokens, medium-resource languages range between 10 million and 1 billion tokens, and low-resource languages fall below 10 million tokens. These languages serve as the training languages in our CPT experiments. The selected training languages, along with their corresponding category and resource level, are summarized in the first three columns of Table 1.
@@ 2.2 Pretraining Data
We finally select data for 15 unique languages, with English appearing in all three categories: altruistic (6 languages, ~92M tokens), selfish (6 languages, ~100M tokens), and stagnant (5 languages, ~87M tokens), resulting in a total of approximately 279 million tokens.
@@
For training configurations that include code, we establish a 2:1 ratio between textual (100 million tokens) and code data (50 million tokens), with code comprising about 33% of the total mixture. We intentionally selected this higher proportion to investigate the effects of a strong code signal within our CPT context. While prior work (Aryabumi et al., 2024) on pretraining from scratch recommends a 25% proportion for balancing language and code performance, it is noted that a higher 33% remains reasonable for enhancing models’ ability, aligning with our goal of studying its trade-offs.
@@ 2.4 CPT Configurations
  • Model $\in\{\text{L3 (Llama-3.1-8B)},\text{L2 (Llama-2-7B)},\text{V7 (Viking-7B)}\}$
  • Data $\in\{\text{Mono (Monolingual)},\text{Bi (Bilingual)}\}$
@@
  • LangCat $\in\{\text{Alt (Altruistic)},\text{Sel (Selfish)},\text{Stag (Stagnant)}\}$
@@ 3.2.2 Comparative Analysis in Classification Tasks
Mid-resource languages show mixed trends. Llama-3.1-8B maintains near-baseline performance with bilingual CPT (64.05% vs baseline 65.85%, -2.7%), but monolingual CPT degrades significantly (52.53%, -20.2%). Llama-2-7B improves accuracy in both configurations: bilingual CPT raises accuracy to 24.26% (vs. baseline 20.59%, +17.8%), while monolingual CPT performs slightly better at 26.80% (+30.2%). Viking-7B uniquely benefits from bilingual CPT (29.74% vs baseline 23.94%, +24.2%), while monolingual CPT underperforms (24.02%, +0.3%). This indicates that bilingual CPT can stabilize mid-resource language performance for certain models (e.g., Viking-7B and Llama-2-7B). However, monolingual CPT risks overfitting to limited in-language data, particularly for models with weaker pretraining (e.g., Llama-3.1-8B).
@@
Low-resource languages exhibit divergent patterns. Llama-3.1-8B improves with bilingual CPT (62.91% vs baseline 56.86%, +10.6%) but declines with monolingual CPT (52.04%, -8.5%). Llama-2-7B improves moderately with bilingual CPT (20.84%, +19.8%) and shows stronger gains with monolingual CPT (24.51%, +40.9%). Viking-7B benefits substantially from bilingual CPT (29.25%, +66.5%), while monolingual CPT slightly underperforms (21.33%, +21.4%). This highlights that bilingual CPT can enhance low-resource language performance for models with compatible pretraining (e.g., Viking-7B and Llama-3.1-8B).
@@ 3.3.1 Downstream Tasks Performance
3.3.2 Language Modeling Performance
To assess the impact of code integration on language modeling, we evaluate performance using Negative Log-Likelihood (NLL) on the MaLA validation set [^7] https://huggingface.co/datasets/MaLA-LM/mala-monolingual-split. We use the implementation of GlotEval (Luo et al., 2025), where the NLL is calculated by concatenating input sentences and using a strided sliding window of 1024 tokens, and is defined as
$NLL=-\sum_{i=1}^{n_{i}}logp_{\theta}(x_{i}|x_{<i}).$
Our results are summarized in Table 2. Across all three base models and all resource levels, the Monolingual+Code configurations yielded lower (better) NLL scores than the monolingual-only CPT. This demonstrates that including code not only helps in downstream classification tasks but also improves the models’ predictive ability and likelihood estimation of unseen text during CPT.
Model High Resource Mid Resource Low Resource
L2-Mono- 5.0080 1.7609 4.3308
L2-Mono+Code- 4.7359 1.7042 4.2289
L3-Mono- 4.8864 2.0375 4.2013
L3-Mono+Code- 4.5701 1.9522 4.1389
V7-Mono- 6.9916 1.9506 4.5367
V7-Mono+Code- 6.6029 1.8369 4.4850
Table 2: NLL across different models and resource levels. Lower is better.
@@ 4 Conclusion
First, while bilingual CPT enhances classification accuracy for mid- and low-resource languages, it introduces language mixing during generation, limiting its utility for translation tasks. Second, code integration during CPT acts as a scaffold for low-resource language understanding but introduces task-dependent trade-offs, improving classification and language modeling capabilities while slightly degrading generation quality. Third, we demonstrate that language classifications based on cross-lingual transfer patterns (altruistic, selfish, stagnant) fail to generalize under varying CPT strategies.
@@
All resources are available at: https://mala-lm.github.io/MixCPT.html.
Acknowledgments
The work has received funding from the Digital Europe Programme under grant agreement No 101195233 (OpenEuroLLM). The authors wish to acknowledge CSC - IT Center for Science, Finland, for providing computational resources.
@@ A.1 Related Work
Continual Pretraining and Language Adaptation
Continual pretraining has emerged as a pivotal technique for adapting LLMs to new domains or languages while retaining previously acquired knowledge (Yıldız et al., 2024). This approach has demonstrated significant benefits across diverse domains, including cybersecurity (Yu et al., 2025), finance(Hirano & Imajo, 2024), and law (Niyogi & Bhattacharya, 2024). In the context of language adaptation, researchers have successfully leveraged continual pretraining to enhance performance on low- and medium-resource languages. For instance, Ji et al. (2024a; 2025); Lu et al. (2024) extended the capabilities of open-weight LLMs by pretraining them on multilingual datasets encompassing hundreds of languages. Similarly, Fujii et al. (2024) significantly improved Japanese language proficiency by continually pretraining LLama-2 (Touvron et al., 2023) on a large-scale Japanese web corpus. Vo et al. (2024) achieved notable advancements in Korean language processing by utilizing 9.7 billion tokens for continual pretraining. Beyond direct continual pretraining, Yong et al. (2023) benchmarks various language adaptation strategies, including adapter-based tuning (MAD-X (Pfeiffer et al., 2020)), and parameter-efficient methods like (IA)³ (Liu et al., 2022), applying them to the BLOOM model under low-resource constraints. Babel (Zhao et al., 2025) introduces a structural expansion of LLMs via layer extension and trains them on curated multilingual corpora, thereby enabling support for 25 widely spoken but underrepresented languages.
@@
Challenges and Dynamics of Cross-Lingual Transfer
While multilingual training can confer benefits, it also introduces the risk of performance degradation, a phenomenon termed the ”curse of multilinguality” or ”negative interference”. Chang et al. (2024) systematically investigated multilingual pretraining and revealed that adding multilingual data can improve performance for low-resource languages up to a certain threshold, beyond which the performance degrades for both low-resource and high-resource languages due to limited model capacity. Moreover, Wang et al. (2020) further demonstrated that negative interference impacts not only high-resource but also low-resource languages, highlighting the universality of this challenge in multilingual modeling. However, multilingual training also presents significant opportunities, particularly through cross-lingual transfer. Protasov et al. (2024) demonstrates that certain high-resource languages, termed ”super donors,” can effectively boost the performance of low-resource ”super recipient” languages in multilingual models.
A.2 CPT Configurations with Structured Naming
Base Model Category Training Data
Mono Bi Mono+Code Bi+Code
Selfish L3-Mono-Sel L3-Bi-Sel L3-Mono+Code-Sel L3-Bi+Code-Sel
Selfish L2-Mono-Sel L2-Bi-Sel L2-Mono+Code-Sel L2-Bi+Code-Sel
Selfish V7-Mono-Sel V7-Bi-Sel V7-Mono+Code-Sel V7-Bi+Code-Sel
A.3 Additional Results on Bilingual CPT with Code Data
@@ A.4 SIB-200 Accuracy
A.5 FLORES-200 BLEU Scores
@@
A.6 Language Mixing Examples